A Complete Bibliography of the Journal of the Royal Statistical Society, Series A family: 1840–1849

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.02

Title word cross-reference

&c [282].

16th [92]. 1841 [94, 92]. 1843 [150, 196]. 1847 [306]. 18th [220].

20th [116, 92]. 25th [220].

30th [174]. 32 [274].

5th [306].

Abstract [45, 94, 102, 268]. Abstracted [146, 145, 139, 142]. Academical [161, 312]. Accidents [311, 94, 217, 111, 95, 142]. According [188]. Account [53, 207, 256, 250, 288, 152, 150, 23, 37, 62, 233]. Accounts [274]. Accurately [166]. Act [66, 274]. Acts [24]. Adaptation [213]. Addition [222]. Additional [319, 234]. Addressed [294, 63]. Adduced [309]. Administration [257]. Admissions [188]. Adopted [323]. Advancement [27, 59, 125, 220, 150, 187, 133]. Affects [83]. Age [188]. Agra [279]. Agricultural [226, 215, 131, 109, 61, 144, 130]. Agriculture [309]. Alleged [242]. America [241, 305]. Among [21, 55, 33, 201, 202, 234, 34, 43, 280]. Amongst [166]. Amount [187]. Analysis [296, 321]. Ancient [53]. Annual [13, 48, 85, 77, 128, 145, 171, 205, 239, 272, 300, 325, 142]. any [301]. Apparatus [178]. Appointed [17, 21, 55, 304, 92].

Aristocracy [234]. Armies [273]. Army [243, 321]. Arrangements [18, 52]. Arranging [35]. Articles [301]. Association [27, 59, 125, 220, 150, 187, 133, 63]. Attached [111]. Attempt [301, 132]. Attended [217]. August [150].

Back [264]. Bank [274]. Banks [231, 124, 216]. Baronetage [201]. be [66, 107]. Bearing [310]. Bedfordshire [144]. been [69]. Before [116, 187]. Being [278, 141, 222]. Belgium [92, 63, 2, 270, 111]. Bengal [198, 267, 321, 318]. Best [17, 166, 35]. Between [301, 155, 7, 245]. Birmingham [15, 5]. Blackrock [183]. Blaenavon [36]. Bolton [83]. Bond [196]. Book [35, 122, 23, 37]. Book-Trade [23]. Borough [119, 281, 56]. both [257]. Boundaries [168, 172]. Bourbons [82, 90]. Braidwood [162]. Brigade [162]. Bristol [65]. Britain [7, 167, 69, 116, 95, 142]. British [187, 27, 59, 125, 220, 150, 63, 133, 129, 257]. Brought [217]. Burman [72].

c [274]. Calculations [241]. Calcutta [197, 321]. Cambridge [220, 100]. cap [66]. Capital [103]. Causes [166, 222]. Census [17, 211, 175, 156, 167, 296, 323, 69]. Central [92, 63]. Cerealia [286]. Certain [160, 175, 71]. Ceylon [4]. Chamber [94]. Chances [96]. Changes [301]. Charges [240]. Charitable [267]. Chaussées [187]. Chiefly [36, 129, 267]. Children [327, 326]. China [245]. Choice [222]. Church [294, 102, 293]. City [199, 191, 160, 33]. Civil [327, 326, 129, 257, 318]. Class [71]. Classes [14, 96, 304, 83, 183, 166, 8, 120, 117, 93]. Cleansing [250]. Climates [210]. Closed [66]. Coal [94, 111]. Collect [21, 55]. Collecting [35]. Collection [221, 235]. Collegiate [102]. Colonel [293]. Colonial [334]. Colonies [280]. Commerce [112]. Commercial [7, 4, 334, 245]. Commission [92, 63, 22]. Commissioners [190, 145, 175, 69]. Committee [17, 21, 55, 89, 162, 138, 119, 180, 304, 133, 293, 14, 306]. Common [35]. Common-Place [35]. Community [96, 166]. Companies [174]. Company [94, 206, 214, 273]. Comparative [96, 71, 167]. Compared [286]. Comparison [202, 43]. Condition [183, 160, 281, 29, 301, 215, 199, 167, 34, 120, 3, 117, 93]. Conducting [167]. Connected [122]. Connexion [301, 155]. Consider [17]. Consisting [293]. Constitution [326, 256]. Consumption [106]. Contained [294]. Continent [73]. Continuation [116, 187]. Contribution [301, 181]. Contributions [131, 137, 151, 178, 221, 235, 161, 312, 5]. Conveyance [159]. Cork [183, 160, 157, 150]. Cornwall [61]. Corporation [53]. Council [17, 13, 48, 85, 77, 128, 171, 278, 304, 306, 293]. Counties [131]. Countries [166, 37, 34, 54]. Country [7]. County [157, 130]. Courts [256]. Crime [139, 251, 302, 268]. Criminal [6, 246, 139, 256, 196, 287, 34, 129, 257]. Currency [225]. Custody [102].

Data [210, 221, 235]. Dated [92]. Dec [282]. December [196, 116]. Decline [2]. Decree [92]. Deduced [233]. Deduction [327]. Denominations [261]. Department [142]. Dependencies [334]. Depression [83]. Derived [167, 221, 235]. Designed [221, 235]. Destitution [110]. Determine [132, 222]. Devastated [141]. Develope [301]. Development [106]. Devon [45]. Devonport [60]. Different [96, 83, 210, 166]. Direction [187]. Disease [43]. Diseases [202]. Dispensaries [267]. Distributed [36]. Distribution [236, 136]. District [278, 157, 322, 24, 36]. Districts [166, 167]. Dividends [282]. Division [136]. Divisions [168, 172]. Dr. [260, 294, 293].

Drawn [187]. Dukinfield [71]. Duration [166, 188, 201, 234, 259, 269]. During [162, 301, 94, 103, 294, 174, 10, 39, 50, 334, 24, 280, 142]. Dustmen [298]. Dwellings [293].

Earliest [23]. East [304, 206, 214, 273]. Eastern [131]. Economical [5]. Economy [155, 216]. Edibles [286]. Education [15, 138, 146, 119, 278, 118, 46, 152, 260, 80, 287, 34, 120, 56, 121, 65]. Educational [331, 277, 313, 328, 330, 206, 214, 113, 130]. Effect [18]. Effects [66, 309]. Eight [301]. Eighth [85, 77]. Eleventh [59, 205]. Elucidate [221, 235]. Empire [72]. Employed [36, 105]. Employment [222]. Employments [137, 151, 181, 222]. End [104]. Ended [94, 139]. Ending [174, 282]. Engine [162]. England [146, 123, 246, 331, 274, 244, 326, 277, 313, 328, 330, 122, 251, 302, 34, 81, 268, 286, 121]. English [78, 88, 234]. Enquire [21, 55]. Enquiry [33, 34]. Enumeration [69, 116]. Erected [107]. Errata [77, 79, 331, 78]. Especially [221, 235]. Esq. [293]. Essays [222]. Establishment [162]. Establishments [73]. Europe [305]. European [21, 55, 273]. Examination [69, 116, 174]. Expenditure [71, 121]. Experience [188]. Extending [301]. Extensive [221, 235]. Extent [168, 172, 231]. Extracts [162].

F [321]. F. [293]. Factory [106]. Facts [241, 35, 310, 309, 69, 116, 287]. Families [201, 71]. Famines [141]. Far [69]. Fatherless [327, 326]. Fellows [304]. Fifteenth [220, 325]. Fifth [138]. Financial [216]. Fine [105]. Finsbury [119]. Fire [162]. Fires [162, 184]. Fisheries [297]. Fluctuations [81]. Following [282]. Food [301]. Former [258]. Forming [210]. Fourteenth [300]. France [152, 80, 309, 187]. Frankfort [191].

Frankfort-on-the-Main [191]. Free [191]. French [152]. Frequent [221, 235]. Friendly [221, 235]. Front [42, 165, 195, 229, 266, 291, 316]. Functions [92]. Funds [327, 326, 62, 108]. Further [70, 10, 151, 161].

G [293]. General [213, 46, 80, 34].

Générale [187]. Genoa [73]. Gentry [234]. George [118, 117, 304]. German [260]. Germany [47, 23]. Giles [293]. Glasgow [29, 143]. Göttingen [260]. Government [168, 172, 267, 321, 318, 121, 279]. Grain [81]. Great [240, 7, 167, 69, 116, 95, 142]. Guardians [107]. Guy [294, 293].

H [293]. Half [282]. Half-Year [282].

Hanover [118, 117]. Harvest [301]. Health [210, 319, 137, 151, 181, 222, 298, 178, 221, 235]. Held [59, 125]. Her [280]. Heritable [136]. High [244]. Hints [215]. Historical [207, 23]. History [173, 124]. Hon. [252]. Horace [252]. Hospital [89, 180]. Hospitals [198]. House [39, 143, 217]. House-Surgeon [217]. Houses [66, 10, 50]. Hull [56, 93]. Hundred [101].

if [301]. II [37]. Illustrating [29].

Illustrations [28, 70, 101, 167]. Improved [157]. Improvement [323]. Improving [215]. Incendiarism [223]. Inclusive [129, 257]. Income [71]. Increase [236, 242, 282, 240]. Increasing [108]. India [206, 214, 273, 141, 129, 206, 214, 257, 286, 267, 273]. Industrial [107, 179]. Industry [187, 270]. Infant [179]. Infirmary [188, 217]. Influence [244, 132, 137, 151, 181, 222, 106, 136, 287]. Influences [221, 235]. Information [1, 12, 32, 41, 68, 76, 99, 115, 149, 164, 186, 194, 219, 228, 255, 265, 285, 290, 308, 315, 333]. Inhabitants [69, 116, 3, 293]. Inhabiting [36]. Inner [118, 117]. Inquiries [167, 120]. Inquiry [213]. Insane [198]. Insanity [189].

Instability [221, 235]. Institutions [78, 88, 206, 214, 111]. Instruction [2, 152]. Intercourse [245]. Investigate [293]. Investigation [304, 301]. Ireland [159, 154, 156, 30, 62, 108, 153]. Irish [211, 175, 158, 297]. Iron [36]. Island [196].

Jails [321]. James [162, 230, 292, 317]. Java [322]. Jews [236]. John [14]. July [59]. June [306, 125, 220, 174]. Jurisdictions [53]. Justice [129, 257, 318]. Juvenile [143].

Kilmurray [157]. Kingdom [17, 301, 334, 139, 201, 174, 231, 124]. Kingston [56, 93]. Kingston-Upon-Hull [56, 93]. Knowledge [137, 151, 181].

Labourers [215]. Labouring [8]. Lancashire [101, 103]. Lane [294, 293]. Large [110]. Last [103, 294, 167]. Law [258, 145]. Laws [221, 235, 190]. Leeds [3]. Letter [18, 252, 294, 63]. Liability [189]. Libraries [47, 305]. Lieut. [293]. Lieut.-Colonel [293]. Life [96, 166, 201, 234, 259, 269]. Lighting [250]. Limerick [33]. List [10, 39, 50]. Literature [122]. Loan [62, 108]. Local [168, 172]. Localities [210]. Locality [221, 235]. London [19, 25, 14, 17, 21, 13, 38, 49, 57, 48, 55, 85, 77, 74, 86, 97, 162, 138, 128, 134, 147, 119, 169, 176, 192, 184, 180, 171, 205, 203, 208, 224, 230, 237, 247, 253, 262, 239, 278, 272, 300, 292, 304, 325, 317, 53, 288, 116, 293]. Lunacy [283, 22]. Lunatic [153]. Lunatics [44].

M. [152]. made [174]. Madras [21, 55, 243]. Main [191]. Maintaining [210, 319]. Majesty [280]. Make [304]. Malton [200]. Management [28, 70, 240]. Managing [162]. Manchester [125, 107, 24, 71, 102, 104, 105]. Mann [252]. Manufacture [158]. Manufacturing [278, 83]. March [92]. Margaret [14]. Markets [288]. Marriage [146, 24]. Marriages [146, 211]. Married [146]. Martyr [16]. Marylebone [188, 120]. Matter [42, 165, 195, 264, 229, 266, 291, 316]. Means [319, 178]. Medical [258, 154, 84]. Meeting [27, 59, 125, 220, 150, 187]. Members [259]. Memoranda [226]. Mendicancy [320]. Merthyr [232]. Method [35, 167]. Metropolis [173, 168, 172, 207, 256, 250]. Metropolitan [66, 162, 22]. Michael [183]. Michaelmas [139]. Middlesex [130]. Military [202, 43]. Mills [105]. Mines [94, 187, 270, 111]. Mining [278, 187, 270]. Minister [152]. Miscellanea [79, 98]. Miscellaneous [11, 20, 26, 31, 40, 51, 58, 67, 75, 87, 114, 127, 135, 148, 163, 170, 177, 185, 193, 204, 209, 218, 227, 238, 248, 254, 263, 271, 276, 284, 289, 299, 303, 307, 314, 324, 329, 332, 335]. MM. [309]. Mode [17, 83]. Modes [166]. Monmouthshire [36]. Monnier [309]. Mont [30]. Months [294]. Monts [73]. Moral [6, 331, 183, 277, 313, 328, 330, 107]. Morning [66]. Mortality [110, 21, 55, 202, 166, 188, 44, 244, 33, 132, 294, 221, 235, 197, 243, 321, 43, 280]. Most [301]. Movement [196]. Mr. [162]. Municipal [78, 88, 250].

Nation [295]. Native [21, 55, 273]. Naval [202, 43]. Navigation [233]. Near [183]. Necessary [301]. Neighbouring [36]. Newcastle [187]. Newspapers [54]. Nice [140]. Nightmen [298]. Ninth [128, 145]. No [146]. Norfolk [226, 196]. North [305, 279, 267]. North-Western [279, 267]. Norway [233]. Notes [190, 61]. Notice [252, 91]. Notices [112]. Nottingham [8]. Nuisances [250]. Number [18, 236]. Numerous [167].

Objects [230, 292, 317]. Observations [319, 326, 234, 175, 122, 156]. Observed [301]. Obtained [178, 69, 116]. Occurred [162]. Occurring [301, 33]. Oct [92]. October [94]. Offenders [6, 34, 143]. Officers [327, 142]. Official [213, 187, 243, 233]. Old [200]. Oldham [94]. Operation [28, 70, 190, 274, 256, 108]. Operations [92]. Origin [102]. Original [221, 235]. Other [188, 73, 34, 286]. Outline [4]. Oxford [100, 249].

P [293, 321]. Paper [146, 116, 187]. Papers [10, 39, 50, 233]. Paris [64, 46, 73]. Parish [45, 118, 144, 102, 117, 183, 157, 36, 120, 16]. Parishes [14, 130]. Parliament [10, 39, 50]. Parochial [258, 145]. Part [36, 37]. Particular [23]. Particularly [101]. Particulars [334]. Parts [160]. Party [318]. Passages [175]. Passengers [159]. Past [101]. Pauper [283]. Pauperism [145, 123]. Paupers [8]. Paving [250]. Pawnbroking [30, 73]. Peerage [201]. People [301, 244]. Pergunnahs [321]. Period [301, 167]. Periodical [141, 91]. Perry [260]. Phenomena [91]. Physical [183]. Piecers [105]. Piete [30, 73]. Place [301, 35]. Places [146, 250, 167]. Plan [35, 323]. Plymouth [59, 60]. Pneumatic [178]. Police [66, 104]. Political [155]. Polytechnic [64]. Ponts [187]. Pontypoll [36]. Poor [28, 70, 258, 145, 107, 33, 153, 190, 145, 283]. Poor-Rates [145]. Poorer [304, 120]. Popular [2]. Population [123, 83, 72, 29, 241, 182, 36, 156, 196, 34, 197]. Populations [166]. Possessed [53]. Postage [18, 52]. Powers [178]. Practical [28, 70]. Practice [84]. Preceding [212]. Prescriptive [53]. Present [10, 101, 83, 173, 187, 231, 207]. Presidency [21, 55, 198]. President [63]. Pressure [166]. Prevailing [202, 294, 43]. Prevalence [242]. Prices [301, 244, 81, 286]. Primary [80]. Principal [47, 261, 305]. Principles [240]. Printed [10, 50, 39].

Prison [139]. Private [84]. Probability [141]. Proceedings [9, 19, 25, 38, 49, 57, 74, 86, 97, 126, 134, 147, 169, 176, 192, 220, 203, 208, 224, 237, 247, 253, 262, 150]. Produce [61, 309]. Produced [240]. Professions [259]. Profitable [240]. Progress [230, 292, 317, 295, 166, 334, 139, 82, 90, 182, 310, 187, 231]. Property [123, 309, 136]. Proposed [35, 167]. Protecting [250]. Provide [327, 326]. Provident [111]. Provinces [236, 279, 267]. Proving [221, 235]. Provisions [250]. Prussia [275, 81]. Prussian [295, 236]. Public [47, 66, 154, 305, 250, 152]. Published [69]. Pulmonary [106]. Punishments [103]. Purposes [213].

Railway [311, 213, 310, 174, 142, 240].

Railways [95, 142, 282]. Rate [283, 221, 235]. Rates [145]. Read [116, 187]. Reading [281, 109]. Real [123, 309]. Receipts [282]. Recently [36, 167]. Recorded [287]. Records [23, 243]. Reductions [240]. Reference [23]. Refuge [143]. Register [94]. Registers [45, 102]. Registration [306, 24]. Regulation [92]. Relation [178]. Relations [7]. Relative [189]. Relief [258, 154]. Religious [261]. Remarks [212, 211, 45, 323]. Reply [212]. Report [190, 258, 14, 17, 21, 15, 13, 64, 48, 55, 85, 77, 89, 162, 138, 128, 145, 119, 180, 171, 205, 239, 278, 272, 300, 304, 306, 325, 183, 160, 188, 84, 44, 175, 46, 152, 24, 80, 179, 3, 56, 133, 293, 321, 142, 93]. Reports [202, 187, 43]. Representing [166]. Requiring [66]. Resources [297]. Respect [188, 174, 34]. Respecting [241]. Respiratory [178]. Results [188, 52, 120]. Returns [225, 211, 166, 139, 213, 174]. Revenue [279, 282]. Review [270]. Rise [245]. Rome [73]. Royal [190, 92]. Rubichon [309]. Rural [157]. Russia [182, 112, 245].

Sabbath [66]. Salaries [327]. Salford [101]. Same [301, 69]. Sanatory [160, 29, 199, 167]. Sanitary [281]. Saving [231, 124, 216]. Scarcities [244]. Scavengers [298]. School [64, 107, 109]. Schools [179]. Science [59, 125, 220, 150, 187, 133]. Sciences [27]. Scotch [70]. Scotland [110, 190, 258]. Scottish [28]. Scrofula [242]. Sea [297]. Seamen [202, 43]. Seasons [132]. sec [66]. Second [55, 180]. Secondary [80]. Secretaries [278]. Section [187, 220, 150, 63]. Self [109]. Self-Supporting [109]. September [27]. Service [326]. Serving [21, 55, 280]. Session [10, 39, 50, 85, 77, 128]. Sessional [146]. Seven [294]. Seventh [48]. Several [259, 130]. Sewerage [173]. Sex [188]. Sexes [189]. Shewn [43]. Showing [83, 282]. Shown [202, 309, 287]. Sicilies [82, 90]. Sicily [7]. Sickness [21, 55, 202, 132, 294, 221, 235, 43]. Silk [158]. since [188]. Sixth [13]. Sixty [103]. Sketch [231]. Social [6]. Societies [9, 221, 235]. Society [19, 25, 14, 17, 21, 13, 38, 49, 57, 48, 55, 85, 77, 74, 86, 97, 138, 128, 134, 147, 119, 169, 176, 192, 180, 171, 205, 203, 208, 224, 230, 237, 247, 253, 262, 239, 278, 272, 300, 292, 304, 325, 317, 116, 293, 126, 306]. Soil [309]. Soldiers [202, 43]. Some [110, 334, 122, 252, 69, 116, 120]. South [296, 278, 159]. Southwark [16]. Sovereigns [269]. Spanish [82, 90]. Spinners [105]. Spinning [105]. Square [230, 292, 317, 118, 117]. St [183, 293, 16]. St. [14, 230, 292, 304, 317, 188, 118, 117]. Staffordshire [278]. State [14, 15, 64, 304, 101, 154, 118, 46, 152, 56, 293, 153]. Statement [294, 71]. States [236, 261, 241, 305, 252]. Statistical [10, 9, 14, 39, 50, 306, 101, 202, 210, 183, 166, 188, 305, 53, 207, 256, 250, 288, 213, 92, 23, 37, 62, 187, 63, 43, 233, 19, 25, 17, 21, 13, 38, 49, 57, 48, 55, 85, 77, 74, 86, 97, 138, 128, 126, 134, 147, 119, 169, 176, 192, 180, 171, 205, 220, 203, 208, 224, 230, 237, 247, 253, 262, 239, 278, 272, 300, 292, 304, 325, 317, 150, 116, 187, 63, 293]. Statistics [47, 21, 6, 55, 89, 184, 180, 283, 331, 83, 281, 140, 226, 320, 157, 4, 200, 29, 322, 326, 8, 78, 88, 131, 173, 277, 313, 328, 330, 275, 100, 249, 260, 178, 252, 36, 232, 155, 196, 144, 221, 235, 251, 302, 161, 312, 34, 104, 105, 54, 5, 22, 129, 191, 206, 214, 198, 257, 279, 267, 273, 318, 113, 130, 16, 60, 65, 268]. Stockport [217]. Stonehouse [60]. Streets [250]. Strong [321]. Subdivision [309]. Suggestions [323]. Sulphur [7]. Summary [138]. Superintendent [162]. Superior [46, 152, 80]. Supplied [221, 235]. Supplying [207]. Supporting [109]. Surgeon [217, 321]. Sussex [109]. System [28, 70, 173, 207, 310, 106, 30].

Table [282]. Tables [211, 246, 217, 287]. Taken [301]. Taking [17, 323]. Tavistock [45]. Ten [294]. Tenth [27, 171]. Territory [241]. Thames [53]. Their [327, 141, 236, 221, 235, 174, 293]. them [210, 250]. Third [181, 142]. Thirteenth [272, 150]. Three [60]. too [221, 235]. Town [93]. Towns [110, 78, 88, 236, 60]. Trade [7, 83, 122, 36, 23, 37, 221, 235, 233]. Traffic [213, 174, 142, 282, 240]. Treatment [188]. Trevethin [36]. Troops [21, 55, 210, 319, 280]. Turin [73]. Tuscany [179]. Twelfth [125, 239]. Twenty [334]. Two [82, 90, 24, 189]. Tydvil [232].

Union [107]. United [17, 261, 241, 301, 334, 305, 139, 201, 252, 174, 231, 124]. Universities [100, 260]. University [249, 260]. Upon [21, 55, 44, 309, 56, 95, 137, 151, 181, 222, 18, 24, 3, 142, 93].

V.P.R.S. [293]. Value [102]. Valuing [178]. Variations [301]. Various [37, 167, 174, 54]. Vic [66]. Vict [274]. View [305, 187].

Villemain [152]. Vital [21, 55, 29, 322, 178, 252, 221, 235, 105, 273]. Volume [1, 12, 32, 41, 68, 76, 99, 115, 149, 164, 186, 194, 219, 228, 255, 265, 285, 290, 308, 315, 333].

W [293]. Wales [146, 123, 246, 331, 277, 313, 328, 330, 251, 296, 302, 34, 268]. Wallingdon [109]. Ward [118, 117]. Water [207]. Wealth [136]. Weather [132]. Week [104]. Werneth [94]. Western [279, 267]. Westminster [14]. Wheat [244]. Which [83, 318, 162, 301, 222]. Widows [327, 326]. Work [260, 309]. Workhouse [107]. Working [14, 183, 24, 71, 117, 93]. Works [36]. Writing [109].

Year [21, 162, 145, 246, 94, 139, 82, 90, 23, 62, 174, 129, 282]. Years [301, 334, 217, 103, 24, 294, 71, 251, 302, 280]. York [199].

References

Anonymous:1840:VIa

[1] Anonymous. Volume information. Journal of the Statistical Society of London, 2(6):467–476, January 1840. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338050.

Rawson:1840:DPI

[2] Rawson W. Rawson. On the decline of popular instruction in Belgium. Journal of the Statistical Society of London, 2(6):385–396, January 1840. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338051.

SCTC:1840:RUC

[3] Statistical Committee of the Town Council. Report upon the condition of the town of Leeds and of its inhabitants.

Journal of the Statistical Society of London, 2(6):397–424, January 1840. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338052.

Capper:1840:OCS

[4] John Capper. An outline of the commercial statistics of Ceylon. Journal of the Statistical Society of London, 2(6):424– 434, January 1840. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338053.

Sub-Committee:1840:CES

[5] Local Sub-Committee. Contributions to the economical statistics of Birmingham. Journal of the Statistical Society of London, 2(6):434–441, January 1840. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338054.

Anonymous:1840:SMS

[6] Anonymous. Social and moral statistics of criminal offenders. Journal of the Statistical Society of London, 2(6):442–445, January 1840. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338055.

Anonymous:1840:STS

[7] Anonymous. On the sulphur trade of Sicily, and the commercial relations between that country and Great Britain. Journal of the Statistical Society of London, 2(6):446–457, January 1840. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338056.

Felkin:1840:SLC

[8] W. Felkin. Statistics of the labouring classes and paupers in Nottingham. Journal of the Statistical Society of London, 2(6):457–459, January 1840. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338057.

Anonymous:1840:PSSa

[9] Anonymous. Proceedings of statistical societies. Journal of the Statistical Society of London, 2(6):459–461, January 1840. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338058.

Anonymous:1840:FLS

[10] Anonymous. A further list of statistical papers printed by the Houses of Parliament during the present session of 1839. Journal of the Statistical Society of London, 2(6):461, January 1840. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338059.

Anonymous:1840:Ma

[11] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 2 (6):462–466, January 1840. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338060.

Anonymous:1840:VIb

[12] Anonymous. Volume information. Journal of the Statistical Society of London, 3 (1):i–iv, April 1840. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337960.

Anonymous:1840:SAR

[13] Anonymous. Sixth annual report of the Council of the Statistical Society of London. Journal of the Statistical Society of London, 3(1):1–13, April 1840. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337961.

Anonymous:1840:RCSa

[14] Anonymous. Report of a Committee of the Statistical Society of London, on the state of the working classes in the parishes of St. Margaret and St. John, Westminster. Journal of the Statistical Society of London, 3(1):14–24, April 1840. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337962.

Anonymous:1840:RSE

[15] Anonymous. Report of the state of education in Birmingham. Journal of the Statistical Society of London, 3(1):25– 49, April 1840. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337963.

Weight:1840:SPS

[16] George Weight. Statistics of the Parish of St. George the Martyr, Southwark. Journal of the Statistical Society of London, 3(1):50–71, April 1840. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337964.

Anonymous:1840:RCSb

[17] Anonymous. Report to the Council of the Statistical Society of London, from the committee appointed to consider the best mode of taking the census of the United Kingdom in 1841. Journal of the Statistical Society of London, 3(1):72– 102, April 1840. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337965.

Hill:1840:ENP

[18] Rowland Hill. On the effect of the new postage arrangements upon the number of letter. Journal of the Statistical Society of London, 3(1):102–105, April 1840. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337966.

Anonymous:1840:PSSb

[19] Anonymous. Proceedings of the Statistical Society of London. Journal of the Statistical Society of London, 3(1):106– 108, April 1840. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337967.

Anonymous:1840:Mb

[20] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 3 (1):109–112, April 1840. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2337968.

Anonymous:1840:RCSc

[21] Anonymous. Report of a committee of the Statistical Society of London, appointed to collect and enquire into vital statistics, upon the sickness and mortality among the European and native troops serving in the Madras presidency, from the year 1793 to 1838. Journal of the Statistical Society of London, 3(2): 113–143, July 1840. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337731.

Sykes:1840:SMC

[22] Lieut.-Col. Sykes. Statistics of the Metropolitan Commission in Lunacy. Journal of the Statistical Society of London, 3(2):143–160, July 1840. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337732.

Meidinger:1840:HSA

[23] M. Henry Meidinger. An historical and statistical account of the book-trade, from the earliest records to the year 1840, with particular reference to Germany. Journal of the Statistical Society of London, 3(2):161–190, July 1840. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2337733.

Johns:1840:RUW

[24] Wm. Johns. Report upon the working of the Registration and Marriage Acts, during the two years 1837–38 and 1838–39, in the Registration District of Manchester. Journal of the Statistical Society of London, 3(2):191–205, July 1840. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2337734.

Anonymous:1840:PSSc

[25] Anonymous. Proceedings of Statistical Society of London. Journal of the Statistical Society of London, 3(2):205– 206, July 1840. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2337735.

Anonymous:1840:Mc

[26] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 3(2): 206–208, July 1840. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2337736.

Anonymous:1840:TMB

[27] Anonymous. Tenth meeting of the British Association for the Advancement of Sciences, September, 1840. Journal of the Statistical Society of London, 3 (3):209–211, October 1840. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337737.

Alison:1840:IPO

[28] W. P. Alison. Illustrations of the practical operation of the Scottish system of management of the poor. Journal of the Statistical Society of London, 3(3):211– 257, October 1840. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337738.

Cowan:1840:VSG

[29] Robert Cowan. Vital statistics of Glasgow, illustrating the sanatory condition of the population. Journal of the Statistical Society of London, 3(3):257–292, October 1840. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337739.

Porter:1840:MPS

[30] Henry John Porter. On the Mont de Piete system of pawnbroking in Ireland. Journal of the Statistical Society of London, 3(3):293–303, October 1840. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2337740.

Anonymous:1840:Md

[31] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 3(3): 304, October 1840. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337741.

Anonymous:1841:VIa

[32] Anonymous. Volume information. Journal of the Statistical Society of London, 3(4):401–408, January 1841. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337717.

Griffin:1841:EMO

[33] Daniel Griffin. An enquiry into the mortality occurring among the poor of the City of Limerick. Journal of the Statistical Society of London, 3(4):305–330, January 1841. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337718.

Rawson:1841:ECC

[34] Rawson W. Rawson. An enquiry into the condition of criminal offenders in England and Wales, with respect to education; or, statistics of education among the criminal and general population of England and other countries. Journal of the Statistical Society of London, 3 (4):331–352, January 1841. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337719.

Guy:1841:BMC

[35] William Augustus Guy. On the best method of collecting and arranging facts, with a proposed new plan of commonplace book. Journal of the Statistical Society of London, 3(4):353–366, January 1841. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337720.

Kenrick:1841:SPP

[36] G. S. Kenrick. Statistics of the population in the Parish of Trevethin (Pontypoll) and at the neighbouring works of Blaenavon in Monmouthshire, chiefly employed in the iron trade, and inhabiting part of the district recently distributed. Journal of the Statistical Society of London, 3(4):366–375, January 1841. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337721.

Meidinger:1841:SAB

[37] M. Henry Meidinger. A statistical account of the book trade of various countries. Part II. Journal of the Statistical Society of London, 3(4):376–386, January 1841. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337722.

Anonymous:1841:PSSa

[38] Anonymous. Proceedings of the Statistical Society of London. Journal of the Statistical Society of London, 3(4):387, January 1841. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337723.

Anonymous:1841:LSPa

[39] Anonymous. A list of statistical papers printed by the house of Parliament, during the session 1840. Journal of the Statistical Society of London, 3(4):387–394, January 1841. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337724.

Anonymous:1841:Ma

[40] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 3 (4):394–400, January 1841. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337725.

Anonymous:1841:VIb

[41] Anonymous. Volume information. Journal of the Statistical Society of London, 4 (1):i–iv, April 1841. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338032.

Anonymous:1841:FM

[42] Anonymous. Front matter. Journal of the Statistical Society of London, 4(1): ii, April 1841. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338033.

Tulloch:1841:CSM

[43] A. M. Tulloch. Comparison of the sickness, mortality, and prevailing disease among seamen and soldiers, as shewn by the Naval and military statistical reports. Journal of the Statistical Society of London, 4(1):1–16, April 1841. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338034.

Farr:1841:RUM

[44] William Farr. Report upon the mortality of lunatics. Journal of the Statistical Society of London, 4(1):17–33, April 1841. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338035.

Barham:1841:RAP

[45] Charles Barham. Remarks on the abstract of the parish registers of Tavistock, Devon. Journal of the Statistical Society of London, 4(1):34–49, April 1841. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338036.

Heywood:1841:RSS

[46] James Heywood. Report on the state of superior general education in Paris. Journal of the Statistical Society of London, 4(1):50–66, April 1841. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338037.

Adrien:1841:SPP

[47] Professor Adrien. Statistics of the principal public libraries in Germany. Journal of the Statistical Society of London, 4(1):66–69, April 1841. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338038.

Anonymous:1841:SAR

[48] Anonymous. Seventh annual report of the Council of the Statistical Society of London. Journal of the Statistical Society of London, 4(1):69–75, April 1841. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338039.

Anonymous:1841:PSSb

[49] Anonymous. Proceedings of the Statistical Society of London. Journal of the Statistical Society of London, 4(1):76– 77, April 1841. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338040.

Anonymous:1841:LSPb

[50] Anonymous. A list of statistical papers printed by the Houses of Parliament, during the session 1841. Journal of the Statistical Society of London, 4(1): 78–79, April 1841. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338041.

Anonymous:1841:Mb

[51] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 4(1): 80–84, April 1841. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338042.

Hill:1841:RNP

[52] Rowland Hill. Results of the new postage arrangements. Journal of the Statistical Society of London, 4(2):85– 99, July 1841. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337884.

Fletcher:1841:SAA

[53] Joseph Fletcher. A statistical account of the ancient prescriptive jurisdictions over the Thames, possessed by the corporation of London. Journal of the Statistical Society of London, 4(2):99–110, July 1841. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2337885.

Simmonds:1841:SNV

[54] P. L. Simmonds. Statistics of newspapers in various countries. Journal of the Statistical Society of London, 4(2):111– 136, July 1841. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337886.

Anonymous:1841:SRC

[55] Anonymous. Second report of a committee of the Statistical Society of London, appointed to collect and enquire into vital statistics, upon the sickness and mortality among the European and native troops serving in the Madras presidency. Journal of the Statistical Society of London, 4(2):137–155, July 1841. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337887.

Society:1841:RSE

[56] Manchester Statistical Society. Report on the state of education in the Borough of Kingston-Upon-Hull. Journal of the Statistical Society of London, 4(2):156– 175, July 1841. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337888.

Anonymous:1841:PSSc

[57] Anonymous. Proceedings of the Statistical Society of London. Journal of the Statistical Society of London, 4(2):175– 176, July 1841. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337889.

Anonymous:1841:Mc

[58] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 4(2): 176–180, July 1841. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337890.

Anonymous:1841:EMB

[59] Anonymous. Eleventh meeting of the British Association for the Advancement of Science. Held at Plymouth, July, 1841. Journal of the Statistical Society of London, 4(3):181–182, October 1841. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979618.

Woolcombe:1841:STT

[60] Henry Woolcombe. Statistics of the three towns of Plymouth, Stonehouse, and Devonport. Journal of the Statistical Society of London, 4(3):183–196, October 1841. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979619.

Lemon:1841:NAP

[61] C. Lemon. Notes on the agricultural produce of Cornwall. Journal of the Statistical Society of London, 4(3):197–208, October 1841. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979620.

Porter:1841:SAL

[62] Henry John Porter. A statistical account of loan funds in Ireland, for the year 1840. Journal of the Statistical Society of London, 4(3):209–224, October 1841. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979621.

Quetelet:1841:LAS

[63] Quetelet and X. Heuschling. Letter addressed to the Statistical Section of the British Association, by the President of the Central Statistical Commission of Belgium. Journal of the Statistical Society of London, 4(3):224–227, October 1841. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979622.

Anonymous:1841:RSP

[64] Anonymous. Report on the state of the Polytechnic School in Paris. Journal of the Statistical Society of London, 4(3):228–249, October 1841. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979623.

ofBristol:1841:SEB

[65] Committee of a Statistical Society of Bristol. Statistics of education in Bristol. Journal of the Statistical Society of London, 4(3):250–267, October 1841. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979624.

Anonymous:1841:EMP

[66] Anonymous. Effects of the Metropolitan Police Act (2 and 3 Vic., cap. 47, sec. 42), requiring public houses to be closed on the Sabbath morning. Journal of the Statistical Society of London, 4(3):268–269, October 1841. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979625.

Anonymous:1841:Md

[67] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 4 (3):269–276, October 1841. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979626.

Anonymous:1842:VIa

[68] Anonymous. Volume information. Journal of the Statistical Society of London, 4(4):363–368, January 1842. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337690.

Porter:1842:ESF

[69] G. R. Porter. An examination of some facts obtained at the recent enumeration of the inhabitants of Great Britain, so far as the same have been published by the Census Commissioners. Journal of the Statistical Society of London, 4(4):277–287, January 1842. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337691.

Alison:1842:FIP

[70] W. P. Alison. Further illustrations of the practical operation of the Scotch system of management of the poor. Journal of the Statistical Society of London, 4(4):288–319, January 1842. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337692.

Neild:1842:CSI

[71] Wm. Neild. Comparative statement of the income and expenditure of certain families of the working class in Manchester and Dukinfield, in the years 1836 and 1841. Journal of the Statistical Society of London, 4(4):320–334, January 1842. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2337693.

Burney:1842:PBE

[72] H. Burney. On the population of the Burman Empire. Journal of the Statistical Society of London, 4(4):335–347, January 1842. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2337694.

Porter:1842:MPR

[73] Henry John Porter. On the Monts de Piete of Rome, Genoa, Turin, and Paris, and other pawnbroking establishments on the Continent. Journal of the Statistical Society of London, 4(4):348–357, January 1842. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2337695.

Anonymous:1842:PSSa

[74] Anonymous. Proceedings of the Statistical Society of London. Journal of the Statistical Society of London, 4(4):358, January 1842. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2337696.

Anonymous:1842:Ma

[75] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 4 (4):359–362, January 1842. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337697.

Anonymous:1842:VIb

[76] Anonymous. Volume information. Journal of the Statistical Society of London, 5 (1):i–iv, April 1842. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2337944.

Anonymous:1842:EEA

[77] Anonymous. Errata: Eighth annual report of the Council of the Statistical Society of London, session 1841–42. Journal of the Statistical Society of London, 5(1):iv, April 1842. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337945. See [85].

Fletcher:1842:ESM

[78] Joseph Fletcher. Errata: Statistics of the municipal institutions of the English towns. Journal of the Statistical Society of London, 5(1):iv, April 1842. CODEN ???? ISSN 0959-5341. URL http:/

/www.jstor.org/stable/2337946. See

[88].

Anonymous:1842:EM

[79] Anonymous. Errata: Miscellanea. Journal of the Statistical Society of London, 5(1):iv, April 1842. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337947.

Jones:1842:RGE

[80] H. L. Jones. Report on general education, superior, secondary, and primary, in France. Journal of the Statistical Society of London, 5(1):1–31, April 1842. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337948.

Rawson:1842:PFG

[81] Rawson W. Rawson. On the prices and fluctuations of grain in Prussia and England, from 1816 to 1841. Journal of the Statistical Society of London, 5(1):32– 46, April 1842. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337949.

Goodwin:1842:PTSa

[82] John Goodwin. Progress of the two Sicilies under the Spanish Bourbons, from the year 1734–35 to 1840. Journal of the Statistical Society of London, 5(1): 47–73, April 1842. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337950.

Ashworth:1842:SPD

[83] Henry Ashworth. Statistics of the present depression of trade at Bolton; showing the mode in which it affects the different classes of a manufacturing population. Journal of the Statistical Society of London, 5(1):74–81, April 1842. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2337951.

Cowan:1842:RPM

[84] Charles Cowan. Report of private medical practice for 1840. Journal of the Statistical Society of London, 5(1):81– 86, April 1842. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337952.

Anonymous:1842:EAR

[85] Anonymous. Eighth annual report of the Council of the Statistical Society of London, session 1841–42. Journal of the Statistical Society of London, 5(1):86– 90, April 1842. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337953. See errata [77].

Anonymous:1842:PSSb

[86] Anonymous. Proceedings of the Statistical Society of London. Journal of the Statistical Society of London, 5(1):91– 92, April 1842. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337954.

Anonymous:1842:Mb

[87] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 5(1): 93–96, April 1842. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337955.

Fletcher:1842:SMI

[88] Joseph Fletcher. Statistics of the municipal institutions of the English towns. Journal of the Statistical Society of London, 5(2):97–168, July 1842. CODEN ???? ISSN 0959-5341. URL http:/

/www.jstor.org/stable/2980708. See

errata [78].

Anonymous:1842:RCH

[89] Anonymous. Report of the committee on hospital statistics. Journal of the Statistical Society of London, 5(2):168– 176, July 1842. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2980709.

Goodwin:1842:PTSb

[90] John Goodwin. Progress of the two Sicilies under the Spanish Bourbons, from the year 1734–35 to 1840. Journal of the Statistical Society of London, 5(2):177– 207, July 1842. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2980710.

Quetelet:1842:NPP

[91] A. Quetelet. Notice on periodical phenomena. Journal of the Statistical Society of London, 5(2):208–209, July 1842. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2980711.

Leopold:1842:RDD

[92] Leopold and King of the Belgians. Royal decree, dated 20th Oct., 1841, for the regulation of the functions and operations of the Central Statistical Commission of Belgium, appointed by a decree of 16th March, 1841. Journal of the Statistical Society of London, 5(2):209– 212, July 1842. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2980712.

ofManchester:1842:RCW

[93] Statistical Society of Manchester. Report on the condition of the working classes in the town of Kingston-upon-Hull. Journal of the Statistical Society of London, 5(2):212–221, July 1842. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2980713.

Fletcher:1842:ARA

[94] Joseph Fletcher. Abstract from a register of accidents in the coal mines of the Chamber and Werneth Company, at Oldham, during the year ended October, 1841. Journal of the Statistical Society of London, 5(2):222–225, July 1842. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2980714.

Weld:1842:AUR

[95] C. R. Weld. On accidents upon the railways in Great Britain. Journal of the Statistical Society of London, 5(2):226– 229, July 1842. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2980715.

Anonymous:1842:CCL

[96] Anonymous. Comparative chances of life in different classes of the community. Journal of the Statistical Society of London, 5(2):230, July 1842. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2980716.

Anonymous:1842:PSSc

[97] Anonymous. Proceedings of the Statistical Society of London. Journal of the Statistical Society of London, 5(2):230– 231, July 1842. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2980717.

Anonymous:1842:Mc

[98] Anonymous. Miscellanea. Journal of the Statistical Society of London, 5(2):231– 234, July 1842. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2980718.

Anonymous:1842:VIc

[99] Anonymous. Volume information. Journal of the Statistical Society of London, 5(3):315–318, October 1842. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337920.

Heywood:1842:SUO

[100] James Heywood and Professor Powell.

Statistics of the universities of Oxford and Cambridge. Journal of the Statistical Society of London, 5(3):235–244, October 1842. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337921.

Ashworth:1842:SIP

[101] Henry Ashworth. Statistical illustrations of the past and present state of Lancashire, and more particularly of the hundred of Salford. Journal of the Statistical Society of London, 5(3):245–256, October 1842. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2337922.

Parkinson:1842:OCV

[102] Richard Parkinson. On the origin, custody, and value of parish registers, with an abstract of the registers of the collegiate Church of Manchester. Journal of the Statistical Society of London, 5(3):256–263, October 1842. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337923.

Hopkins:1842:CPL

[103] R. W. Hopkins. Capital punishments in Lancashire during the last sixty years. Journal of the Statistical Society of London, 5(3):263–265, October 1842. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2337924.

Shaw:1842:PSW

[104] Charles Shaw. Police statistics of the Week’s end in Manchester. Journal of the Statistical Society of London, 5(3): 266–268, October 1842. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2337925.

Shuttleworth:1842:VSS

[105] John Shuttleworth. Vital statistics of the spinners and piecers employed in the fine spinning Mills of Manchester. Journal of the Statistical Society of London, 5(3):268–273, October 1842. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337926.

Noble:1842:IFS

[106] D. Noble. On the influence of the factory system in the development of pulmonary consumption. Journal of the Statistical Society of London, 5(3):274– 280, October 1842. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337927.

Gardiner:1842:MIW

[107] Ner Gardiner. On the moral and industrial workhouse school about to be erected by the guardians of the poor of the Manchester Union. Journal of the Statistical Society of London, 5(3):280– 282, October 1842. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337928.

Porter:1842:IOL

[108] Henry John Porter. On the increasing operation of loan funds in Ireland. Journal of the Statistical Society of London, 5(3):282–288, October 1842. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337929.

Gilbert:1842:SSR

[109] Davies Gilbert. On the self-supporting reading, writing, and agricultural school, at Wallingdon, in Sussex. Journal of the Statistical Society of London, 5(3):289, October 1842. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337930.

Alison:1842:DMS

[110] W. P. Alison. On the destitution and mortality in some of the large towns in Scotland. Journal of the Statistical Society of London, 5(3):289–292, October 1842. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337931.

Weld:1842:ACM

[111] C. R. Weld. On accidents in coal mines in Belgium and the provident institutions attached to the mines. Journal of the Statistical Society of London, 5 (3):292–300, October 1842. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337932.

Slowaczynski:1842:NCR

[112] A. Slowaczynski. Notices of the commerce of Russia. Journal of the Statistical Society of London, 5(3):300–306, October 1842. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2337933.

Tremenheere:1842:ES

[113] Seymour Tremenheere. Educational statistics, 1840–41. Journal of the Statistical Society of London, 5(3):306–308, October 1842. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2337934.

Anonymous:1842:Md

[114] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 5 (3):308–314, October 1842. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337935.

Anonymous:1843:VIa

[115] Anonymous. Volume information. Journal of the Statistical Society of London, 6 (1):i–iv, April 1843. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338006.

Porter:1843:ESF

[116] G. R. Porter. An examination of some facts obtained at the recent enumeration of the inhabitants of Great Britain, in continuation of a paper read before the Statistical Society of London, on the 20th December, 1841. Journal of the Statistical Society of London, 6(1): 1–16, April 1843. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338007.

Weld:1843:CWC

[117] C. R. Weld. On the condition of the working classes in the inner ward of St. George’s Parish, Hanover Square. Journal of the Statistical Society of London, 6(1):17–23, April 1843. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338008.

Edgell:1843:SEI

[118] E. Wyatt Edgell. On the state of education in the inner ward of St. George’s Parish, Hanover Square. Journal of the Statistical Society of London, 6(1):24– 27, April 1843. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338009.

Anonymous:1843:REC

[119] Anonymous. Report of the education committee of the Statistical Society of London on the Borough of Finsbury. Journal of the Statistical Society of London, 6(1):28–43, April 1843. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338010.

Rawson:1843:RSI

[120] R. W. Rawson. Results of some inquiries into the condition and education of the poorer classes in the Parish of Marylebone, in 1838. Journal of the Statistical Society of London, 6(1):44–48, April 1843. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338011.

Tremenheere:1843:GEE

[121] Seymour Tremenheere. Government expenditure on education in England, 1841–42. Journal of the Statistical Society of London, 6(1):48–50, April 1843. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338012.

James:1843:SOB

[122] G. P. R. James. Some observations on the book trade, as connected with literature, in England. Journal of the Statistical Society of London, 6(1):50–60, April 1843. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338013.

Anonymous:1843:RPP

[123] Anonymous. Real property, population, and pauperism in England and Wales. Journal of the Statistical Society of London, 6(1):60–72, April 1843. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338014.

Pratt:1843:HSB

[124] John Tidd Pratt. The history of savings’ banks in the United Kingdom. Journal of the Statistical Society of London, 6(1): 73–74, April 1843. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338015.

Anonymous:1843:TMB

[125] Anonymous. Twelfth meeting of the British Association for the Advancement of Science. Held at Manchester, June, 1842. Journal of the Statistical Society of London, 6(1):75–76, April 1843. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338016.

Anonymous:1843:PSSa

[126] Anonymous. Proceedings of the Statistical Society. Journal of the Statistical Society of London, 6(1):76, April 1843. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338017.

Anonymous:1843:Ma

[127] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 6(1): 77–88, April 1843. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338018.

Anonymous:1843:NAR

[128] Anonymous. Ninth annual report of the Council of the Statistical Society of London. Session, 1842–3. Journal of the Statistical Society of London, 6(2):89– 93, May 1843. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337865.

Sykes:1843:SCC

[129] W. H. Sykes. Statistics of civil and criminal justice in British India, chiefly from the year 1836 to 1840 inclusive. Journal of the Statistical Society of London, 6(2): 94–119, May 1843. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337866.

Tremenheere:1843:AES

[130] Henry Tremenheere. Agricultural and educational statistics of several parishes in the County of Middlesex. Journal of the Statistical Society of London, 6(2): 120–130, May 1843. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337867.

Fletcher:1843:CAS

[131] Joseph Fletcher. Contributions to the agricultural statistics of the Eastern counties. Journal of the Statistical Society of London, 6(2):130–133, May 1843. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337868.

Guy:1843:ADI

[132] William Augustus Guy and M. B. Cantab. An attempt to determine the influence of the seasons and weather on sickness and mortality. Journal of the Statistical Society of London, 6(2):133– 150, May 1843. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337869.

Sykes:1843:RCB

[133] W. H. Sykes, Lord Sandon, G. R. Porter, J. Heywood, and W. P. Alison. Report of a committee of the British Association for the Advancement of Science. Journal of the Statistical Society of London, 6(2):150–166, May 1843. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337870.

Anonymous:1843:PSSb

[134] Anonymous. Proceedings of the Statistical Society of London. Journal of the Statistical Society of London, 6(2):166– 168, May 1843. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337871.

Anonymous:1843:Mb

[135] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 6(2): 168–184, May 1843. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337872.

Passy:1843:DHP

[136] M. Passy and John P. Boileau. On the division of heritable property, and its influence on the distribution of wealth. Journal of the Statistical Society of London, 6(3):185–197, August 1843. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2337788.

Guy:1843:CKI

[137] William Augustus Guy. Contributions to a knowledge of the influence of employments upon health. Journal of the Statistical Society of London, 6(3):197– 211, August 1843. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337789.

Anonymous:1843:FRS

[138] Anonymous. Fifth report and summary of the education committee of the Statistical Society of London. Journal of the Statistical Society of London, 6(3):211– 218, August 1843. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337790.

Fletcher:1843:PCU

[139] Joseph Fletcher. Progress of crime in the United Kingdom: Abstracted from the criminal returns for 1842, and the prison returns for the year ended at Michaelmas, 1841. Journal of the Statistical Society of London, 6(3):218–240, August 1843. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337791.

Boileau:1843:SN

[140] John P. Boileau. Statistics of nice. Journal of the Statistical Society of London, 6(3):240–246, August 1843. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337792.

Everest:1843:FDI

[141] Robert Everest. On the famines that have devastated India, and on the probability of their being periodical. Journal of the Statistical Society of London, 6(3):246–248, August 1843. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337793.

Weld:1843:ATU

[142] C. R. Weld. Accidents and traffic upon the railways in Great Britain, during 1842; abstracted from the third annual report of the Officers of the Railway Department. Journal of the Statistical Society of London, 6(3):249–252, August 1843. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2337794.

Tancred:1843:HRJ

[143] T. Tancred. House of refuge for juvenile offenders, Glasgow. Journal of the Statistical Society of London, 6(3):252– 255, August 1843. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2337795.

Martin:1843:SAP

[144] John Edward Martin. Statistics of an agricultural parish in Bedfordshire. Journal of the Statistical Society of London, 6(3):255, August 1843. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337796.

Anonymous:1843:PPR

[145] Anonymous. Pauperism and poor-rates in the parochial year 1842; abstracted from the ninth annual report of the Poor Law Commissioners, 1843. Journal of the Statistical Society of London, 6(3): 256–259, August 1843. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2337797.

Anonymous:1843:PMM

[146] Anonymous. Places of marriage, marriages, and education of the married, in England and Wales; abstracted from the sessional paper, no. 223, 1843. Journal of the Statistical Society of London, 6(3):260–261, August 1843. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337798.

Anonymous:1843:PSSc

[147] Anonymous. Proceedings of the Statistical Society of London. Journal of the Statistical Society of London, 6(3): 261, August 1843. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337799.

Anonymous:1843:Mc

[148] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 6(3): 261–280, August 1843. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2337800.

Anonymous:1843:VIb

[149] Anonymous. Volume information. Journal of the Statistical Society of London, 6(4):377–388, December 1843. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337891.

Heywood:1843:TMB

[150] James Heywood. Thirteenth meeting of the British Association for the Advancement of Science, at Cork, August 17–23, 1843: Account of the proceedings of the Statistical Section. Journal of the Statistical Society of London, 6(4):281–282, December 1843. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2337892.

Guy:1843:FCK

[151] William Augustus Guy. Further contributions to a knowledge of the influence of employments upon health. Journal of the Statistical Society of London, 6 (4):283–304, December 1843. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337893.

Heywood:1843:ARM

[152] James Heywood. Account of a report of M. Villemain, the French Minister of Public Instruction, on the state of superior education in France. Journal of the Statistical Society of London, 6 (4):304–310, December 1843. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337894.

Taylor:1843:SLP

[153] W. Cooke Taylor. State of the lunatic poor in Ireland. Journal of the Statistical Society of London, 6(4):311–317, December 1843. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2337895.

Bullen:1843:SPM

[154] Dr. Bullen. State of public medical relief in Ireland. Journal of the Statistical Society of London, 6(4):317–322, December 1843. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2337896.

Lawson:1843:CBS

[155] Professor Lawson. On the connexion between statistics and political economy. Journal of the Statistical Society of London, 6(4):322, December 1843. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337897.

Larcom:1843:OCP

[156] Captain Larcom. Observations on the census of the population of Ireland in 1841. Journal of the Statistical Society of London, 6(4):323–351, December 1843. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337898.

Bullen:1843:SIR

[157] Dr D. Bullen. Statistics of an improved rural district (the Parish of Kilmurray) in the County of Cork. Journal of the Statistical Society of London, 6(4):352–354, December 1843. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337899.

Taylor:1843:ISM

[158] W. Cooke Taylor. On the Irish silk manufacture. Journal of the Statistical Society of London, 6(4):354–356, December 1843. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337900.

Bianconi:1843:CPS

[159] Mr. Bianconi. On conveyance of passengers in the South Ireland. Journal of the Statistical Society of London, 6 (4):356–357, December 1843. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337901.

Biggs:1843:RSC

[160] Henry Biggs. Report on the sanatory condition of certain parts of the City of Cork. Journal of the Statistical Society of London, 6(4):357–360, December 1843. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337902.

Powell:1843:FCA

[161] Rev. Professor Powell. Further contributions to academical statistics. Journal of the Statistical Society of London, 6(4):360, December 1843. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337903.

Anonymous:1843:ERS

[162] Anonymous. Extracts from the report of the Superintendent of the Metropolitan Fire Brigade (Mr. James Braidwood) to the committee for managing the London fire engine establishment, on the fires which occurred during the year 1842. Journal of the Statistical Society of London, 6(4):361–362, December 1843. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2337904.

Anonymous:1843:Md

[163] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 6 (4):363–376, December 1843. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337905.

Anonymous:1844:VIa

[164] Anonymous. Volume information. Journal of the Statistical Society of London, 7 (1):i–iv, April 1844. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337742.

Anonymous:1844:FM

[165] Anonymous. Front matter. Journal of the Statistical Society of London, 7(1): ii, April 1844. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337743.

Chadwick:1844:BMR

[166] Edwin Chadwick. On the best modes of representing accurately, by statistical returns, the duration of life, and the pressure and progress of the causes of mortality amongst different classes of the community, and amongst the populations of different districts and countries. Journal of the Statistical Society of London, 7(1):1–40, April 1844. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337744.

Neison:1844:MRP

[167] F. G. P. Neison. On a method recently proposed for conducting inquiries into the comparative sanatory condition of various districts, with illustrations, derived from numerous places in Great Britain at the period of the last census. Journal of the Statistical Society of London, 7(1):40–68, April 1844. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337745.

Fletcher:1844:MBEa

[168] Joseph Fletcher. The Metropolis: Its boundaries, extent, and divisions for local government. Journal of the Statistical Society of London, 7(1):69–85, April 1844. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337746.

Anonymous:1844:PSSa

[169] Anonymous. Proceedings of the Statistical Society of London. Journal of the Statistical Society of London, 7(1):86– 87, April 1844. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2337747.

Anonymous:1844:Ma

[170] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 7(1): 87–96, April 1844. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2337748.

Anonymous:1844:TAR

[171] Anonymous. Tenth annual report of the Council of the Statistical Society of London. Journal of the Statistical Society of London, 7(2):97–103, June 1844. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338086.

Fletcher:1844:MBEb

[172] Joseph Fletcher. The Metropolis; its boundaries, extent, and divisions for local government. Journal of the Statistical Society of London, 7(2):103–143, June 1844. CODEN ???? ISSN 0959-

                                        5341. URL http://www.jstor.org/
                                        stable/2338087.

Fletcher:1844:HSP

[173] Joseph Fletcher. History and statistics of the present system of sewerage in the Metropolis. Journal of the Statistical Society of London, 7(2):143–170, June 1844. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338088.

Porter:1844:ERM

[174] G. R. Porter. An examination of the returns made by the various railway companies of the United Kingdom, with respect to their traffic during the year ending 30th June, 1843. Journal of the Statistical Society of London, 7(2):170– 178, June 1844. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338089.

Hallam:1844:OCP

[175] Henry Hallam. Observations on certain passages in the report of the Irish Census Commissioners. Journal of the Statistical Society of London, 7(2):178– 180, June 1844. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338090.

Anonymous:1844:PSSb

[176] Anonymous. Proceedings of the Statistical Society of London. Journal of the Statistical Society of London, 7(2):181– 182, June 1844. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338091.

Anonymous:1844:Mb

[177] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 7(2): 182–192, June 1844. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338092.

Hutchinson:1844:CVS

[178] John Hutchinson. Contributions to vital statistics, obtained by means of a pneumatic apparatus for valuing the respiratory powers with relation to health. Journal of the Statistical Society of London, 7(3):193–212, September 1844. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2337936.

Mayer:1844:RII

[179] Enrico Mayer. Report on the infant industrial schools of Tuscany. Journal of the Statistical Society of London, 7 (3):213–214, September 1844. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337937.

Anonymous:1844:SRC

[180] Anonymous. Second report of the committee of the Statistical Society of London on hospital statistics. Journal of the Statistical Society of London, 7(3): 214–231, September 1844. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2337938.

Guy:1844:TCK

[181] William Augustus Guy. A third contribution to a knowledge of the influence of employments upon health. Journal of the Statistical Society of London, 7 (3):232–243, September 1844. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337939.

Graham:1844:PPR

[182] Major Graham. On the progress of the population of Russia. Journal of the Statistical Society of London, 7(3):243–250, September 1844. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2337940.

Beamish:1844:SRP

[183] North Ludlow Beamish. Statistical report on the physical and moral condition of the working classes in the Parish of St. Michael, Blackrock, near Cork. Journal of the Statistical Society of London, 7(3):251–254, September 1844. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337941.

Anonymous:1844:SFL

[184] Anonymous. Statistics of fires in London. Journal of the Statistical Society of London, 7(3):255–265, September 1844. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337942.

Anonymous:1844:Mc

[185] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 7 (3):266–280, September 1844. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337943.

Anonymous:1844:VIb

[186] Anonymous. Volume information. Journal of the Statistical Society of London, 7(4):369–371, December 1844. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337969.

Porter:1844:SVR

[187] G. R. Porter. A statistical view of the recent progress and present amount of mining industry in France; drawn from the official reports of the “Direction Générale des Ponts et Chaussées et des Mines,” in continuation of a paper read before the Statistical Section of the British Association for the Advancement of Science, at its meeting in Newcastle in 1838. Journal of the Statistical Society of London, 7(4):281–291, December 1844. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337970.

Clendinning:1844:RES

[188] John Clendinning. Report on the experience of the St. Marylebone Infirmary, since 1827, with respect to admissions, duration of treatment, mortality, and other statistical results, according to age and sex. Journal of the Statistical Society of London, 7(4):292–310, December 1844. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2337971.

Thurnam:1844:RLT

[189] John Thurnam. On the relative liability of the two sexes to insanity. Journal of the Statistical Society of London, 7(4):310–316, December 1844. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337972.

Alison:1844:NRR

[190] J. P. Alison. Notes on the report of the Royal Commissioners on the operation of the Poor Laws in Scotland, 1844. Journal of the Statistical Society of London, 7(4):316–318, December 1844. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2337973.

Sykes:1844:SFC

[191] W. H. Sykes. The statistics of the Free City of Frankfort-on-the-Main. Journal of the Statistical Society of London, 7 (4):318–360, December 1844. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337974.

Anonymous:1844:PSSc

[192] Anonymous. Proceedings of the Statistical Society of London. Journal of the Statistical Society of London, 7(4):360, December 1844. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337975.

Anonymous:1844:Md

[193] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 7 (4):361–368, December 1844. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337976.

Anonymous:1845:VIa

[194] Anonymous. Volume information. Journal of the Statistical Society of London, 8(1):i–iv, March 1845. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337991.

Anonymous:1845:FM

[195] Anonymous. Front matter. Journal of the Statistical Society of London, 8(1): ii, March 1845. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337992.

Maconochie:1845:CSM

[196] Captain Maconochie. Criminal statistics and movement of the bond population of Norfolk Island, to December, 1843. Journal of the Statistical Society of London, 8(1):1–49, March 1845. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337993.

Sykes:1845:PMC

[197] W. H. Sykes. On the population and mortality of Calcutta. Journal of the Statistical Society of London, 8(1):50– 58, March 1845. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337994.

Sykes:1845:SHI

[198] W. H. Sykes. Statistics of the hospitals for the insane under the Bengal presidency. Journal of the Statistical Society of London, 8(1):58–63, March 1845. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2337995.

Laycock:1845:SCC

[199] T. Laycock. On the sanatory condition of the city of York. Journal of the Statistical Society of London, 8(1):63– 66, March 1845. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337996.

Copperthwaite:1845:SON

[200] William Charles Copperthwaite. Statistics of Old and New Malton. Journal of the Statistical Society of London, 8(1): 66–68, March 1845. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337997.

Guy:1845:DLA

[201] William A. Guy. On the duration of life among the families of the Peerage and Baronetage of the United Kingdom. Journal of the Statistical Society of London, 8(1):69–77, March 1845. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337998.

Balfour:1845:CSM

[202] T. Graham Balfour. Comparison of the sickness, mortality, and prevailing diseases among seamen and soldiers, as shown by the Naval and military statistical reports. Journal of the Statistical Society of London, 8(1):77–86, March 1845. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337999.

Anonymous:1845:PSSa

[203] Anonymous. Proceedings of the Statistical Society of London. Journal of the Statistical Society of London, 8(1): 86, March 1845. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338000.

Anonymous:1845:Ma

[204] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 8(1): 87–96, March 1845. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338001.

Anonymous:1845:EAR

[205] Anonymous. Eleventh annual report of the Statistical Society of London. Journal of the Statistical Society of London, 8(2):97–103, June 1845. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337726.

Sykes:1845:SEIa

[206] W. H. Sykes. Statistics of the educational institutions of the East India Company in India. Journal of the Statistical Society of London, 8(2):103–147, June 1845. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2337727.

Fletcher:1845:HSA

[207] Joseph Fletcher. Historical and statistical account of the present system of supplying the Metropolis with water. Journal of the Statistical Society of London, 8(2):148–181, June 1845. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337728.

Anonymous:1845:PSSb

[208] Anonymous. Proceedings of the Statistical Society of London. Journal of the Statistical Society of London, 8(2): 182, June 1845. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337729.

Anonymous:1845:Mb

[209] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 8(2): 182–192, June 1845. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337730.

Balfour:1845:SDF

[210] Edward Balfour. Statistical data for forming troops and maintaining them in health in different climates and localities. Journal of the Statistical Society of London, 8(3):193–209, September 1845. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337839.

Anonymous:1845:RTM

[211] Anonymous. Remarks on tables of marriages in the Irish census returns for 1841. Journal of the Statistical Society of London, 8(3):209–214, September 1845. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337840.

Anonymous:1845:RPR

[212] Anonymous. Reply to the preceding remarks. Journal of the Statistical Society of London, 8(3):214–215, September 1845. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337841.

Graham:1845:AOR

[213] W. A. Graham. Adaptation of official returns of railway traffic to the general purposes of statistical inquiry. Journal of the Statistical Society of London, 8 (3):215–236, September 1845. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337842.

Sykes:1845:SEIb

[214] W. H. Sykes. Statistics of the educational institutions of the East India Company in India. Journal of the Statistical Society of London, 8(3):236–273, September 1845. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337843.

Dury:1845:HIC

[215] Theodore Dury. Hints for improving the condition of agricultural labourers. Journal of the Statistical Society of London, 8(3):273–275, September 1845. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2337844.

Woollgar:1845:FES

[216] J. W. Woollgar. On the financial economy of savings’ banks. Journal of the Statistical Society of London, 8(3):275– 277, September 1845. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2337845.

Gaskell:1845:TAB

[217] Samuel Gaskell. Tables of accidents brought to the Stockport Infirmary, and attended by the House-Surgeon, in the years 1833, 1834 and 1835. Journal of the Statistical Society of London, 8 (3):277–281, September 1845. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337846.

Anonymous:1845:Mc

[218] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 8 (3):281–288, September 1845. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337847.

Anonymous:1845:VIb

[219] Anonymous. Volume information. Journal of the Statistical Society of London, 8(4):369–372, December 1845. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337779.

Anonymous:1845:FMB

[220] Anonymous. Fifteenth meeting of the British Association for the Advancement of Science, at Cambridge, June 18th– 25th, 1845. proceedings of the Statistical Section. Journal of the Statistical Society of London, 8(4):289–290, December 1845. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2337780.

Neison:1845:CVS

[221] F. G. P. Neison. Contributions to vital statistics, especially designed to elucidate the rate of mortality, the laws of sickness, and the influences of trade and locality on health, derived from an extensive collection of original data, supplied by friendly societies, and proving their too frequent instability. Journal of the Statistical Society of London, 8 (4):290–350, December 1845. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337781.

Guy:1845:CWD

[222] William Augustus Guy. On the causes which determine the choice of an employment; being an addition to the essays on the influence of employments upon health. Journal of the Statistical Society of London, 8(4):351–353, December 1845. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337782.

Anonymous:1845:I

[223] Anonymous. Incendiarism. Journal of the Statistical Society of London, 8(4): 353–354, December 1845. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2337783.

Anonymous:1845:PSSc

[224] Anonymous. Proceedings of the Statistical Society of London. Journal of the Statistical Society of London, 8(4): 354–358, December 1845. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2337784.

Anonymous:1845:CR

[225] Anonymous. Currency returns. Journal of the Statistical Society of London, 8(4):358–359, December 1845. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337785.

Boileau:1845:MTA

[226] John P. Boileau. Memoranda towards the agricultural statistics of Norfolk. Journal of the Statistical Society of London, 8(4):360, December 1845. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337786.

Anonymous:1845:Md

[227] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 8 (4):361–368, December 1845. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337787.

Anonymous:1846:VIa

[228] Anonymous. Volume information. Journal of the Statistical Society of London, 9(1):i–iv, March 1846. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337873.

Anonymous:1846:FM

[229] Anonymous. Front matter. Journal of the Statistical Society of London, 9(1): ii, March 1846. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2337874.

Anonymous:1846:OPS

[230] Anonymous. Objects and progress of the Statistical Society of London, 12, St. James’s Square. Journal of the Statistical Society of London, 9(1):1–2, March 1846. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2337875.

Porter:1846:SPP

[231] G. R. Porter. Sketch of the progress and present extent of savings’ banks in the United Kingdom. Journal of the Statistical Society of London, 9(1):1–13, March 1846. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2337876.

Kenrick:1846:SMT

[232] G. S. Kenrick. Statistics of Merthyr Tydvil. Journal of the Statistical Society of London, 9(1):14–21, March 1846. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2337877.

Valpy:1846:SAT

[233] Richard Valpy. Statistical account of the trade and navigation of Norway, deduced from official papers. Journal of the Statistical Society of London, 9(1): 22–36, March 1846. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337878.

Guy:1846:DLA

[234] William A. Guy. On the duration of life among the English gentry, with additional observations on the duration of life among the aristocracy. Journal of the Statistical Society of London, 9(1): 37–49, March 1846. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337879.

Neison:1846:CVS

[235] F. G. P. Neison. Contributions to vital statistics, especially designed to elucidate the rate of mortality, the laws of sickness, and the influences of trade and locality on health, derived from an extensive collection of original data, supplied by friendly societies, and proving their too frequent instability. Journal of the Statistical Society of London, 9(1): 50–76, March 1846. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337880.

Hoffman:1846:NIJ

[236] M. Hoffman and C. R. Weld. On the number and increase of the Jews in the Prussian states, and their distribution in the provinces and towns. Journal of the Statistical Society of London, 9(1):77– 85, March 1846. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337881.

Anonymous:1846:PSSa

[237] Anonymous. Proceedings of the Statistical Society of London. Journal of the Statistical Society of London, 9(1): 85, March 1846. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2337882.

Anonymous:1846:Ma

[238] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 9(1): 85–96, March 1846. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2337883.

Anonymous:1846:TAR

[239] Anonymous. Twelfth annual report of the Statistical Society of London. Journal of the Statistical Society of London, 9(2):97–100, June 1846. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337829.

Williams:1846:PRM

[240] J. Butler Williams. On the principles of railway management, and on the profitable increase in the traffic produced by great reductions in the charges. Journal of the Statistical Society of London, 9(2): 101–132, June 1846. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2337830.

Currie:1846:FCR

[241] Dr. Currie. Facts and calculations respecting the population and territory of the United States of America. Journal of the Statistical Society of London, 9(2): 133–151, June 1846. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2337831.

Phillips:1846:PAI

[242] Benjamin Phillips. The prevalence and alleged increase of scrofula. Journal of the Statistical Society of London, 9(2): 152–157, June 1846. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337832.

Sykes:1846:MMA

[243] Lieut.-Colonel Sykes. Mortality of the Madras Army; from official records. Journal of the Statistical Society of London, 9(2):157, June 1846. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2337833.

Farr:1846:ISH

[244] William Farr. The influence of scarcities and of the high prices of wheat on the mortality of the people of England. Journal of the Statistical Society of London, 9(2):158–174, June 1846. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337834.

Valpy:1846:RCI

[245] Richard Valpy. The rise of commercial intercourse between Russia and China. Journal of the Statistical Society of London, 9(2):175–176, June 1846. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337835.

Anonymous:1846:CTY

[246] Anonymous. Criminal tables for the year 1845. — England and Wales. Journal of the Statistical Society of London, 9(2): 177–183, June 1846. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337836.

Anonymous:1846:PSSb

[247] Anonymous. Proceedings of the Statistical Society of London. Journal of the Statistical Society of London, 9(2): 184, June 1846. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337837.

Anonymous:1846:Mb

[248] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 9(2): 184–192, June 1846. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337838.

Heywood:1846:OUS

[249] James Heywood. Oxford University statistics. Journal of the Statistical Society of London, 9(3):193–203, October 1846. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337859.

Fletcher:1846:SAM

[250] Joseph Fletcher. A statistical account of the municipal provisions for paving, lighting, and cleansing the streets and public places of the Metropolis, and for protecting them from nuisances. Journal of the Statistical Society of London, 9(3):204–222, October 1846. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337860.

Neison:1846:SCE

[251] F. G. P. Neison. Statistics of crime in England and Wales for the years 1842, 1843, and 1844. Journal of the Statistical Society of London, 9(3):223–276, October 1846. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337861.

Jarvis:1846:NSV

[252] Edward Jarvis and Thos Laycock. Notice of some vital statistics of the United States, in a letter to the Hon. Horace Mann. Journal of the Statistical Society of London, 9(3):277–279, October 1846. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337862.

Anonymous:1846:PSSc

[253] Anonymous. Proceedings of the Statistical Society of London. Journal of the Statistical Society of London, 9(3):280, October 1846. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337863.

Anonymous:1846:Mc

[254] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 9 (3):280–288, October 1846. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337864.

Anonymous:1846:VIb

[255] Anonymous. Volume information. Journal of the Statistical Society of London, 9(4):377–380, December 1846. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337675.

Fletcher:1846:SAC

[256] Joseph Fletcher. Statistical account of the constitution and operation of the criminal courts of the Metropolis. Journal of the Statistical Society of London, 9(4):289–309, December 1846. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337676.

Sykes:1846:SAC

[257] W. H. Sykes. Statistics of the administration of civil and criminal justice in British India from 1841 to 1844, both inclusive. Journal of the Statistical Society of London, 9(4):310–338, December 1846. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337677.

Alison:1846:RMR

[258] Dr. Alison. Report on the medical relief to the parochial poor of Scotland, under the former poor law. Journal of the Statistical Society of London, 9(4):339–345, December 1846. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337678.

Guy:1846:DLM

[259] William A. Guy. On the duration of life in the members of the several professions. Journal of the Statistical Society of London, 9(4):346–353, December 1846. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337679.

Heywood:1846:SGU

[260] James Heywood. Statistics of the German universities, from a recent work on German University education by Dr. Perry, of Göttingen. Journal of the Statistical Society of London, 9(4):354–364, December 1846. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337680.

Anonymous:1846:PRD

[261] Anonymous. Principal religious denominations in the United States. Journal of the Statistical Society of London, 9(4):365, December 1846. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337681.

Anonymous:1846:PSSd

[262] Anonymous. Proceedings of the Statistical Society of London. Journal of the Statistical Society of London, 9(4):365, December 1846. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337682.

Anonymous:1846:Md

[263] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 9 (4):366–376, December 1846. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337683.

Anonymous:1846:BM

[264] Anonymous. Back matter. Journal of the Statistical Society of London, 9(4):1– 32, December 1846. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337684.

Anonymous:1847:VIa

[265] Anonymous. Volume information. Journal of the Statistical Society of London, 10(1):i–iv, March 1847. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338025.

Anonymous:1847:FM

[266] Anonymous. Front matter. Journal of the Statistical Society of London, 10(1): ii, March 1847. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338026.

Sykes:1847:SGC

[267] W. H. Sykes. Statistics of the government charitable dispensaries of India, chiefly in the Bengal and North-Western Provinces. Journal of the Statistical Society of London, 10(1):1–37, March 1847. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338027.

Russell:1847:ASC

[268] Whitworth Russell. Abstract of the “Statistics of crime in England and Wales, from 1839 to 1843”. Journal of the Statistical Society of London, 10(1): 38–61, March 1847. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338028.

Guy:1847:DLS

[269] William A. Guy. On the duration of life of sovereigns. Journal of the Statistical Society of London, 10(1):62–69, March 1847. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338029.

Valpy:1847:RMM

[270] Richard Valpy. A review of the mines and mining industry of Belgium. Journal of the Statistical Society of London, 10(1):70–83, March 1847. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338030.

Anonymous:1847:Ma

[271] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 10(1): 84–96, March 1847. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338031.

Anonymous:1847:TAR

[272] Anonymous. Thirteenth annual report of the Statistical Society of London. Journal of the Statistical Society of London, 10(2):97–99, May 1847. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337685.

Sykes:1847:VSE

[273] W. H. Sykes. Vital statistics of the East India Company’s armies in India, European and native. Journal of the Statistical Society of London, 10(2):100– 131, May 1847. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337686.

Danson:1847:ABE

[274] J. T. Danson. On the accounts of the Bank of England under the operation of the Act 7 & 8 Vict., c. 32. Journal of the Statistical Society of London, 10(2):132– 153, May 1847. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337687.

Hebeler:1847:SP

[275] Bernard Hebeler. Statistics of Prussia.

Journal of the Statistical Society of London, 10(2):154–186, May 1847. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337688.

Anonymous:1847:Mb

[276] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 10(2): 187–192, May 1847. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337689.

Fletcher:1847:MES

[277] Joseph Fletcher. Moral and educational statistics of England and Wales. Journal of the Statistical Society of London, 10 (3):193–233, September 1847. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337698.

Anonymous:1847:EMM

[278] Anonymous. Education in the mining and manufacturing district of South Staffordshire; being a report to the Council of the Statistical Society of London by its Secretaries. Journal of the Statistical Society of London, 10(3):234– 242, September 1847. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2337699.

Sykes:1847:RSA

[279] W. H. Sykes. Revenue statistics of the Agra Government, or North-Western Provinces. Journal of the Statistical Society of London, 10(3):243–251, September 1847. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2337700.

Tulloch:1847:MAH

[280] A. M. Tulloch. On the mortality among Her Majesty’s troops serving in the Colonies during the years 1844 and 1845. Journal of the Statistical Society of London, 10(3):252–259, September 1847. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2337701.

Billing:1847:SSC

[281] John Billing. Statistics of the sanitary condition of the Borough of reading. Journal of the Statistical Society of London, 10(3):259–261, September 1847. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2337702.

Whishaw:1847:TSR

[282] J. Whishaw. A table showing the revenue receipts, increase of traffic, dividends, &c., on the following railways for the half-year ending in dec., 1846. Journal of the Statistical Society of London, 10(3):262–266, September 1847. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2337703.

Anonymous:1847:PLP

[283] Anonymous. Pauper lunacy and Poor’s rate statistics. Journal of the Statistical Society of London, 10(3):266–267, September 1847. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337704.

Anonymous:1847:Mc

[284] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 10 (3):268–288, September 1847. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337705.

Anonymous:1847:VIb

[285] Anonymous. Volume information. Journal of the Statistical Society of London, 10(4):377–379, November 1847. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337812.

Sykes:1847:PCO

[286] W. H. Sykes. Prices of the cerealia and other edibles in India and England compared. Journal of the Statistical Society of London, 10(4):289–316, November 1847. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337813.

Porter:1847:IES

[287] G. R. Porter. The influence of education, shown by facts recorded in the criminal tables for 1845 and 1846. Journal of the Statistical Society of London, 10 (4):316–344, November 1847. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337814.

Fletcher:1847:SAM

[288] Joseph Fletcher. Statistical account of the markets of London. Journal of the Statistical Society of London, 10(4):345– 360, November 1847. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2337815.

Anonymous:1847:Md

[289] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 10 (4):361–376, November 1847. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337816.

Anonymous:1848:VIa

[290] Anonymous. Volume information. Journal of the Statistical Society of London, 11(1):i–iv, March 1848. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337759.

Anonymous:1848:FM

[291] Anonymous. Front matter. Journal of the Statistical Society of London, 11(1): ii, March 1848. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2337760.

Anonymous:1848:OPS

[292] Anonymous. Objects and progress of the Statistical Society of London, 12, St. James’s Square. Journal of the Statistical Society of London, 11(1):1–2, March 1848. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2337761.

Sykes:1848:RCC

[293] W. H. Sykes and F. G. P. Neison. Report of a committee of the Council of the Statistical Society of London, consisting of Lieut.-Colonel W. H. Sykes, V.P.R.S., Dr. Guy, and F. G. P. Neison, Esq., to investigate the state of the inhabitants and their dwellings in Church Lane, St. Giles’s. Journal of the Statistical Society of London, 11(1):1–18, March 1848. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2337762.

Mann:1848:SMP

[294] Horace Mann. Statement of the mortality prevailing in Church lane during the last ten years, with the sickness during the last seven months. contained in a letter addressed to Dr. Guy. Journal of the Statistical Society of London, 11(1):19– 24, March 1848. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337763.

Banfield:1848:PPN

[295] T. C. Banfield. The progress of the Prussian nation, 1805, 1831, 1842. Journal of the Statistical Society of London, 11(1): 25–37, March 1848. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337764.

Neison:1848:ACN

[296] F. G. P. Neison. Analysis of the census of New South Wales. Journal of the Statistical Society of London, 11(1):38– 54, March 1848. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337765.

Valpy:1848:RIS

[297] Richard Valpy. The resources of the Irish sea fisheries. Journal of the Statistical Society of London, 11(1):55–72, March 1848. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337766.

Guy:1848:HNS

[298] William Augustus Guy. On the health of nightmen, scavengers, and dustmen. Journal of the Statistical Society of London, 11(1):72–81, March 1848. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337767.

Anonymous:1848:Ma

[299] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 11(1): 81–96, March 1848. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337768.

Anonymous:1848:FAR

[300] Anonymous. Fourteenth annual report of the Statistical Society of London. Journal of the Statistical Society of London, 11(2):97–100, May 1848. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338002.

Danson:1848:CTI

[301] J. T. Danson. A contribution towards an investigation of the changes which have taken place in the condition of the people of the United Kingdom during the eight years extending from the harvest of 1839 to the harvest of 1847; and an attempt to develope the connexion (if any), between the changes observed and the variations occurring during the same period in the prices of the most necessary articles of food. Journal of the Statistical Society of London, 11(2):101– 140, May 1848. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338003.

Neison:1848:SCE

[302] F. G. P. Neison. Statistics of crime in England and Wales, for the years 1834– 1844. Journal of the Statistical Society of London, 11(2):140–165, May 1848. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338004.

Anonymous:1848:Mb

[303] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 11(2): 166–192, May 1848. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338005.

Anonymous:1848:RCS

[304] Anonymous. Report to the Council of the Statistical Society of London from a committee of its Fellows appointed to make an investigation into the state of the poorer classes in St. George’s in the East. Journal of the Statistical Society of London, 11(3):193–249, August 1848. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337956.

Edwards:1848:SVP

[305] Edward Edwards. A statistical view of the principal public libraries in Europe and the United States of North America. Journal of the Statistical Society of London, 11(3):250–281, August 1848. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2337957.

Anonymous:1848:RRC

[306] Anonymous. Report of the Registration Committee to the Council of the Statistical Society, 5th June, 1847. Journal of the Statistical Society of London, 11(3):282–287, August 1848. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337958.

Anonymous:1848:Mc

[307] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 11(3): 288–304, August 1848. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2337959.

Anonymous:1848:VIb

[308] Anonymous. Volume information. Journal of the Statistical Society of London, 11(4):381–384, November 1848. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338043.

Lovelace:1848:SRP

[309] Earl Lovelace. On the subdivision of real property and its effects upon agriculture and the produce of the soil in France, shown by the facts adduced in the recent work of MM. Monnier and Rubichon. Journal of the Statistical Society of London, 11(4):305–322, November 1848. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338044.

Harding:1848:FBP

[310] Wyndham Harding. Facts bearing on the progress of the railway system. Journal of the Statistical Society of London, 11(4):322–343, November 1848. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338045.

Anonymous:1848:RA

[311] Anonymous. Railway accidents. Journal of the Statistical Society of London, 11(4):343, November 1848. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338046.

Powell:1848:CAS

[312] Professor Powell. Contributions to academical statistics. Journal of the Statistical Society of London, 11(4):344, November 1848. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338047.

Fletcher:1848:MES

[313] Joseph Fletcher. Moral and educational statistics of England and Wales. Journal of the Statistical Society of London, 11 (4):344–366, November 1848. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338048.

Anonymous:1848:Md

[314] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 11 (4):367–380, November 1848. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338049.

Anonymous:1849:VIa

[315] Anonymous. Volume information. Journal of the Statistical Society of London, 12(1):i–iv, February 1849. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338241.

Anonymous:1849:FM

[316] Anonymous. Front matter. Journal of the Statistical Society of London, 12(1): ii, February 1849. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338242.

Anonymous:1849:OPS

[317] Anonymous. Objects and progress of the Statistical Society of London 12, St. James’s Square. Journal of the Statistical Society of London, 12(1):1–2, February 1849. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338243.

Sykes:1849:SCJ

[318] Colonel Sykes. The statistics of civil justice in Bengal, in which the government is a party. Journal of the Statistical Society of London, 12(1):3–33, February 1849. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338244.

Balfour:1849:AOM

[319] Edward Balfour. Additional observations on the means of maintaining troops in health. Journal of the Statistical Society of London, 12(1):33–42, February 1849. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338245.

Boileau:1849:SM

[320] John P. Boileau. Statistics of mendicancy. Journal of the Statistical Society of London, 12(1):43–48, February 1849. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338246.

Sykes:1849:ARS

[321] W. H. Sykes. Analysis of the report of Surgeon F. P. Strong, of the Bengal Army, to the Bengal government for 1847, of the mortality in the jails of the 24 Pergunnahs, Calcutta. Journal of the Statistical Society of London, 12(1):48– 59, February 1849. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338247.

Crawfurd:1849:VSD

[322] John Crawfurd and W. H. Sykes. Vital statistics of a district in Java. Journal of the Statistical Society of London, 12(1):60–71, February 1849. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338248.

Wyatt-Edgell:1849:RPA

[323] E. Wyatt-Edgell. Remarks on the plan adopted for taking the census in 1841, with suggestions for its improvement. Journal of the Statistical Society of London, 12(1):71–77, February 1849. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338249.

Anonymous:1849:Ma

[324] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 12(1): 77–96, February 1849. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338250.

Anonymous:1849:FAR

[325] Anonymous. Fifteenth annual report of the Statistical Society of London. Journal of the Statistical Society of London, 12(2):97–102, May 1849. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338219.

Farr:1849:SCS

[326] William Farr. Statistics of the civil service of England, with observations on the constitution of funds, to provide for fatherless children and widows. Journal of the Statistical Society of London, 12(2):103–124, May 1849. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338220.

Anonymous:1849:FPF

[327] Anonymous. On funds to provide for the fatherless children and widows of civil officers–or of others–from a deduction out of their salaries. Journal of the Statistical Society of London, 12(2):125– 150, May 1849. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338221.

Fletcher:1849:MESa

[328] Joseph Fletcher. Moral and educational statistics of England and Wales. Journal of the Statistical Society of London, 12(2):151–176, May 1849. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338222.

Anonymous:1849:Mb

[329] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 12(2): 177–188, May 1849. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338223.

Fletcher:1849:MESb

[330] Joseph Fletcher. Moral and educational statistics of England and Wales. Journal of the Statistical Society of London, 12(3):189–335, August 1849. CODEN ???? ISSN 0959-5341. URL http:/

/www.jstor.org/stable/2338176. See

errata [331].

Anonymous:1849:EME

[331] Anonymous. Errata: Moral and educational statistics of England and Wales. Journal of the Statistical Society of London, 12(3):335, August 1849. CODEN ???? ISSN 0959-5341. URL http:/

/www.jstor.org/stable/2338177. See

[330].

Anonymous:1849:Mc

[332] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 12(3):

336–348, August 1849. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338178.

Anonymous:1849:VIb

[333] Anonymous. Volume information. Journal of the Statistical Society of London, 12(4):449–452, November 1849. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338275.

Danson:1849:SPC

[334] J. T. Danson. Some particulars of the commercial progress of the colonial dependencies of the United Kingdom, during the twenty years, 1827–46. Journal of the Statistical Society of London, 12 (4):349–439, November 1849. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338276.

Anonymous:1849:Md

[335] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 12 (4):439–448, November 1849. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338277.